OpenStudy (anonymous):

Evaluate the expression. 4 2

OpenStudy (anonymous):

Mhm?

OpenStudy (anonymous):

There's just a 4 and a 2, I'm not sure what you want us to do here :o

OpenStudy (anonymous):

(4/2) evluate the expression

OpenStudy (anonymous):

oh 4/2 = 2

OpenStudy (anonymous):

I think that's what you're saying o_O

ganeshie8 (ganeshie8):

4 c 2 or 4 p 2

ganeshie8 (ganeshie8):

may be.. @iambatman

ganeshie8 (ganeshie8):

could u take screenshot of question u seeing and attach ha ? :)

OpenStudy (anonymous):

^

OpenStudy (anonymous):

ok hold on

ganeshie8 (ganeshie8):

\(\large \mathbb{\binom{4}{2}} \)

ganeshie8 (ganeshie8):

its read as \(\large ^4C_2\)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

wow are good

OpenStudy (anonymous):

@ganeshie8 He's the best! :P

OpenStudy (anonymous):

I see

ganeshie8 (ganeshie8):

\(\large ^nC_r = \frac{n!}{(n-r)!r!}\)

ganeshie8 (ganeshie8):

\(\large ^4C_2 = \frac{4!}{(4-2)!2!}\)

ganeshie8 (ganeshie8):

u familiar wid factorials right ?

ganeshie8 (ganeshie8):

knw how to evaluate \(4!\) ?

OpenStudy (anonymous):

naw

ganeshie8 (ganeshie8):

to take factorial of a number, simply multiply it wid all the natural numbers before that number : \(4! = 4 \times 3 \times 2 \times 1\)

ganeshie8 (ganeshie8):

\(2! = 2\times 1\)

OpenStudy (anonymous):

math is my weakest subject

ganeshie8 (ganeshie8):

Next, see if u can simplify below : \(\large ^4C_2 = \frac{4!}{(4-2)!2!}\) \(\large~~ ~~~~= ?\)

ganeshie8 (ganeshie8):

you're doing great :)

ganeshie8 (ganeshie8):

\(\large ^4C_2 = \frac{4!}{(4-2)!2!}\) \(\large ^4C_2 = \frac{4!}{2!2!}\) \(\large ^4C_2 = \frac{4\times 3 \times 2 \times 1}{2\times 1\times 2 \times 1}\) \(\large ^4C_2 = \frac{24}{4}\) \(\large ^4C_2 = 6\)

ganeshie8 (ganeshie8):

see if that makes more or less sense...

OpenStudy (anonymous):

so do you see how many ways it can be multiplied

ganeshie8 (ganeshie8):

kindof yes